Please help me solve this:
Find the equation of the line through (3,4) and (6,8)

Answer by ewatrrr(24785)

 
Hi,
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
Find the equation of the line through
(3,4) and
(6,8) m = -4/-3 = 4/3
y = (4/3)x + b ||using Pt (3,4) to solve fo b
4 = 4 + b
0 = b
y = (4/3)x